OpenStudy (samanthagreer):

What is the value of the expression when a = 7, b = 10, and c = 2?

Mehek (mehek14):

So substitute the values

OpenStudy (dtan5457):

\[\frac{ 8(7) }{ 10+2(2) }\]

OpenStudy (dtan5457):

56/14=?

Mehek (mehek14):

8*7/10+2*2 56/10+4 56/14 4

OpenStudy (samanthagreer):

omg thank i @Mehek14 i tried and tried and couded figure it out thank u

OpenStudy (danjs):

use parenthesis mehek, the way you wrote it in the first line comes to \[8\frac{ 7 }{ 10 } + 2*2\]

Mehek (mehek14):

ok @DanJS no problem @samanthagreer